A coffee shop has 4 different types of coffee. One can order his/her coffee in a

Small, Medium, or Large cup. He/she must also choose to add at least one of

the following; Gream, Sugar, or Milk (any combination is possible). In how many

ways can you order a coffee?


1
Expert's answer
2022-02-01T10:23:31-0500

 You have four options for coffee, three options for size, two options for cream, two options for sugar, and two options for milk, and you have to pick one option for each of these five choices.

"4\cdot 3\cdot 2^3=96" ways